# GraphTeam: Facilitating Large Language Model-based Graph Analysis via Multi-Agent Collaboration

### v2

@@ Abstract.
Graphs are widely used for modeling relational data in real-world scenarios, such as social networks and urban computing. While large language models (LLMs) have achieved strong performance in many areas, existing LLM-based graph analysis approaches either integrate graph neural networks (GNNs) for specific machine learning tasks (e.g., node classification), limiting their transferability, or rely solely on LLMs’ internal reasoning ability, resulting in suboptimal performance. To address these limitations, we take advantage of recent advances in LLM-based agents, which have shown capabilities of utilizing external knowledge or tools for problem solving. By simulating human problem-solving strategies such as analogy and collaboration, we propose a multi-agent system based on LLMs named GraphTeam, for graph analysis. GraphTeam consists of five LLM-based agents from three modules, and the agents with different specialities can collaborate with each other to address complex problems. Specifically, (1) input-output normalization module: the question agent extracts and refines four key arguments (e.g., graph type and output format) from the original question, facilitating the problem understanding, and the answer agent organizes the results to meet the output requirement; (2) external knowledge retrieval module: we first build a knowledge base consisting of relevant documentation and experience information, and then the search agent retrieves the most relevant entries from the knowledge base for each question. (3) problem-solving module: given the retrieved information from search agent, the coding agent uses established algorithms via programming to generate solutions, and in case the coding agent does not work, the reasoning agent will directly compute the results without programming. Extensive experiments on six graph analysis benchmarks demonstrate that GraphTeam achieves state-of-the-art performance with an average 25.85% improvement over the best baseline in terms of accuracy. The code and data are available at https://github.com/BUPT-GAMMA/GraphTeam.
Large Language Model, Multi-agent System, Graph Analysis
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Graph is a commonly used data structure for modeling relational data. Correspondingly, tasks and methods of graph analysis have been extensively exploited in AI field (Brown et al., 2020; Team et al., 2024; Anthropic, 2024) and achieved significantly success in many fields, such as social network (Myers et al., 2014; Min et al., 2021) and unban computing (Yan et al., 2024; Zou et al., 2024). Recently, with the surge of LLM, some researchers begin to explore the feasibility of graph analysis with LLMs (Tang et al., 2024; Zhang et al., 2024a; Ouyang et al., 2024; Sanford et al., 2024).
Existing LLM-based graph analysis methods mainly fall into two categories. The first category (Chai et al., 2023; Liu et al., 2024, 2023; Qin et al., 2023) usually combines LLMs with graph neural networks (GNNs) for modeling. A typical paradigm is to first encode the input graph with a GNN, and then feed the node encodings to an LLM as language token embeddings. However, this line of work is specialized for graph machine learning tasks, especially node classification, and has limited transferability to general graph analysis problems. The second category (Fatemi et al., 2023; Chen et al., 2024b) usually flattens graphs into textual descriptions, and purely relies on LLMs for analyzing (Zhao et al., 2023; Shi et al., 2023; Huang et al., 2023; Liu and Wu, 2023; Chen et al., 2024c). Strategies like in-context learning (Brown et al., 2020) and chain-of-thought reasoning (Wei et al., 2023) are widely used in these methods to improve performance. However, these methods analyze graphs by step-by-step reasoning on raw inputs without utilizing any external knowledge or tools, severely limiting their performance.
Fortunately, LLMs have been shown to be capable of using external knowledge or tools for solving problems, known as a key feature of LLM-based agents (Xi et al., 2023; Wang et al., 2024b). Moreover, inspired by the division of labor and cooperation mechanisms in human society, researchers propose LLM-based multi-agent systems (Chen et al., 2023; Qian et al., 2024; Hong et al., 2023; Wu et al., 2023), where agents with different specialities can discuss and collaborate with each other to address complex problems. However, previous multi-agent frameworks are designed for general purposes and have few knowledge relevant to graph analysis, and thus fail to outperform the aforementioned methods tailored for graph analysis.
@@ 1. Introduction
Figure 1. The overall pipeline of our multi-agent system GraphTeam (left), and the comparison between GraphTeam and state-of-the-art baseline on six benchmarks (right).
@@
To overcome the limitations of previous work, we aim to design a multi-agent system based on LLMs with its design and implementation specialized for graph analysis. Inspired by the problem-solving process of humans (Robson and Pólya, 1945; Ramnarine-Rieks, 2011; Bybee et al., 2006; Brand-Gruwel et al., 2009; Savery and Duffy, 1995; Eisenberg and Berkowitz, 1990), we outline our system GraphTeam with three core modules: an input-output normalization module, an external knowledge retrieval module, and a problem-solving module. The overall pipeline of GraphTeam is shown in Fig. 1. (1) During human problem-solving processes, clearly understanding and formalizing the problem is the first step (Bybee et al., 2006; Robson and Pólya, 1945), and correctly answering the question in required formats is also important. Therefore, we introduce the input-output normalization module with two different LLM-based agents: the question agent aims to extract and refine four key arguments (e.g., graph type and output format) from the original question description, facilitating the system to better understand the problem; the answer agent aims to reformat the results to meet the output requirement. (2) During the solving process of a hard problem, humans will seek help from external sources (Eisenberg and Berkowitz, 1990) or make analogy to previously solved ones (Bybee et al., 2006). Therefore, we first build a knowledge base consisting of relevant documentation and experience information. Then we introduce the external knowledge retrieval module with a search agent to retrieve the most relevant entries from the knowledge base for each question. The retrieved information is expected to benefit the downstream solving process. (3) In the problem-solving module, we introduce a coding agent that uses established algorithms via programming to generate solutions. If the coding agent fails to generate effective codes, a reasoning agent will directly infer the results without programming. We conduct extensive experiments on six graph analysis benchmarks, including Talk like a Graph (Fatemi et al., 2023), LLM4DyG (Zhang et al., 2024b), GraphWiz (Chen et al., 2024a), NLGraph (Wang et al., 2024a), Graph Instruct (Luo et al., 2024) and a newly introduced GNN-AutoGL based on AutoGL (Guan et al., 2021). These benchmarks cover a variety of graph tasks, ranging from basic graph theory to the deployment of GNNs. The experimental results and ablation studies demonstrate the effectiveness of GraphTeam as well as the necessity of each module. In summary, our contributions are three-fold:
@@
$\bullet$ Different from previous paradigms that utilize LLMs for graph analysis, we propose to facilitate the analysis via the collaboration of multiple LLM-based agents.
@@
$\bullet$ Inspired by the problem-solving process of humans, we propose a carefully designed framework GraphTeam with five agents from three functional groups, i.e., input-output normalization, external knowledge retrieval, and problem-solving modules. The agents can work in cooperation with a due division of labor.
@@
$\bullet$ We conduct extensive experiments over very recent graph analysis benchmarks. Experiments show that GraphTeam achieves state-of-the-art (SOTA) performance on all six benchmarks, with an average 25.85% improvement on accuracy over the best baseline.
2. Related Work
2.1. LLM-based Graph Analysis
2.1.1. Benchmarks
@@
To evaluate the ability of large language models (LLMs) in solving graph analysis problems, researchers have developed various benchmarks (Das et al., 2024; Chen et al., 2024a; Guo et al., 2023; Fatemi et al., 2023). Among them, NLGraph (Wang et al., 2024a) is one of the first benchmark for assessing LLMs’ graph analysis capabilities, focusing on basic graph theory reasoning and simple GNN computations. Graphwiz (Chen et al., 2024a) evaluates LLMs on graph problems of varying complexity, including NP-complete problems. GraphInstruct (Luo et al., 2024) includes 21 graph analysis tasks covering node-level, node-pair-level, and graph-level problems. Talk like a Graph (Fatemi et al., 2023) measures basic graph understanding abilities, while LLM4DyG (Zhang et al., 2024b) focuses on dynamic graph tasks related to temporal and spatial information. In our experiments, we employ the above five benchmarks and also build an additional one focusing on GNNs for assessing the abilities of different methods.
2.1.2. Methods
@@
There is a rising trend of research to utilize LLMs for various graph analysis tasks (Li et al., 2024b; Jin et al., 2024). Here we focus on the most relevant methods that employ LLMs as predictors (Chai et al., 2023; Liu et al., 2024, 2023; Qin et al., 2023). Related work can be roughly divided into two categories. The first category typically integrates LLMs with graph neural network (GNNs), and is specialized for graph machine learning tasks such as node classification (Tang et al., 2024; Zhang et al., 2024a). For example, GraphGPT (Tang et al., 2024) employs instruction tuning to fine-tune a GNN-enhanced LLM for node classification. GraphTranslator (Zhang et al., 2024a) encodes nodes in a graph by a GNN, and then project node embeddings into the embedding space of LLM tokens for zero-shot classification. However, these methods are designed and optimized for a specific task, and cannot generalize well to the diverse analysis tasks in the above benchmarks. The second category usually flatten graphs into textual descriptions of adjacency lists, and then rely on LLM itself to analyze graphs with strategies like in-context learning (Brown et al., 2020), chain-of-thought reasoning (Wei et al., 2023) for inference (Zhao et al., 2023; Shi et al., 2023; Huang et al., 2023; Liu and Wu, 2023; Chen et al., 2024c). For example, GUNDAM (Ouyang et al., 2024) collects natural language-based question-answer pairs of graph analysis problems to fine-tune LLMs, and performs step-by-step reasoning for inference. However, recent studies have shown that the reasoning depths of current LLMs are still shallow (Lam et al., 2024; Li et al., 2024a), and will probably fail for graph analysis problems that require complex reasoning. In contrast, we encourage LLMs to leverage external knowledge and tools, instead of performing step-by-step reasoning on raw inputs.
2.2. Multi-Agent Collaboration of LLMs
Autonomous agents aim to solve problems with self-directed planning and actions. Recently, researchers introduce the concept of LLM-based agent that is capable of understanding natural language instruction, interacting with humans, perceiving external environments, and performing various actions (Xi et al., 2023; Bang et al., 2023; Li et al., 2023; Wang et al., 2024b). Moreover, inspired by the collaboration of human in solving problems, researchers introduce LLM-based multi-agent systems, where multiple agents work in cooperation with a due division of labor.
@@ 2.2. Multi-Agent Collaboration of LLMs
For example, ChatDev (Qian et al., 2024) is a virtual software company where LLM-based agents are responsible for different software development tasks (such as user interface design or software testing) for collaboration. MetaGPT (Hong et al., 2023) is a meta-programming framework that simulates human workflows. With proper developing pipeline, MetaGPT can generate effective programming solutions. AgentVerse (Chen et al., 2023) is a multi-agent framework inspired by human collaboration dynamics. It involves four iterative stages of expert recruitment, decision discussion, action and evaluation. AutoGen (Wu et al., 2023) is another multi-agent framework for general purposes. Through tool invocation and code generation, AutoGen can accomplish a wide variety of tasks. MACRec (Wang et al., 2024c) is a novel framework designed to enhance recommendation systems through multi-agent collaboration. This system includes various specialized agents and can easily solve various recommendation tasks. However, these multi-agent frameworks are not specialized for graph analysis, e.g., they have few knowledge of relevant Python libraries and no experience base of similar problems. Consequently, they can hardly beat the SOTA performance on the aforementioned graph analysis benchmarks.
3. Methodology
@@
In this section, we propose GraphTeam, a multi-agent system based on LLMs for graph analysis. We will start by introducing the overall framework, and then present the details of each agent in the system.
3.1. Framework Overview
As shown in Fig. 2, the entire framework of GraphTeam consists of five agents from three functional groups, i.e., question and answer agents for input-output normalization, search agent for external knowledge retrieval, as well as coding and reasoning agents for problem solving.
@@ 3.1. Framework Overview
Figure 2. The overall framework of GraphTeam, which includes five agents from three functional groups.
@@
External Knowledge Retrieval: This functional group only includes the search agent to extract relevant information from external knowledge base. Specifically, we first build a knowledge base for graph analysis based on documentation of popular Python libraries and experiences of previous problem-solving processes. With problems refined by the upstream question agent as queries, the search agent then retrieves information beneficial to downstream problem solving from the knowledge base.
Problem Solving: This functional group aims to calculate the results based on the refined problem of question agent and the extracted knowledge of search agent. Specifically, the coding agent attempts to write Python codes to solve the problem. When the generated codes cannot be executed normally, a retry mechanism will be employed to fix the errors in the codes. If the coding agent still fails after several retries, the reasoning agent will solve the problem directly without programming.
3.2. Question Agent
Abstraction is a preliminary step for human beings to solve problems (Robson and Pólya, 1945). Inspired by this, we introduce a question agent to extract and refine four key arguments from the original problem descriptions. Formally, given a graph analysis problem with its original description denoted as $q$, we employ an LLM denoted as $\operatorname{Question}(q)$ to generate refined question $q_{r}$, graph type $q_{t}$, input graph $q_{g}$ and output format $q_{f}$:
(1) $(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q),$
where $q_{r}$ is a condensed version of original instruction, $q_{t}$ indicates the graph type (e.g., directed, undirected, weighted, etc.), $q_{g}$ contains the graph to be analyzed, and $q_{f}$ specifies how the answer should be presented (e.g., numerical value, list, etc.). By explicitly identifying these key information, downstream agents can better understand a task and its requirements, thereby enhancing the overall analysis capability. An example of the question agent is shown below:

### v3

# GraphTeam: Facilitating Large Language Model-based Graph Reasoning via Multi-Agent Collaboration

@@ Abstract
Graphs are widely used for modeling relational data in real-world scenarios, such as social networks and urban computing. While large language models (LLMs) have achieved strong performance in many areas, existing LLM-based graph analysis approaches either integrate graph neural networks (GNNs) for specific machine learning tasks (e.g., node classification), limiting their transferability, or rely solely on LLMs’ internal reasoning ability, resulting in suboptimal performance. To address these limitations, we take advantage of recent advances in LLM-based agents, which have shown capabilities of utilizing external knowledge or tools for problem solving. By simulating human problem-solving strategies such as analogy and collaboration, we propose a multi-agent system based on LLMs named GraphTeam, for graph analysis. GraphTeam consists of five LLM-based agents from three modules, and the agents with different specialities can collaborate with each other to address complex problems. Specifically, (1) input-output normalization module: the question agent extracts and refines four key arguments (e.g., graph type and output format) from the original question, facilitating the problem understanding, and the answer agent organizes the results to meet the output requirement; (2) external knowledge retrieval module: we first build a knowledge base consisting of relevant documentation and experience information, and then the search agent retrieves the most relevant entries from the knowledge base for each question. (3) problem-solving module: given the retrieved information from search agent, the coding agent uses established algorithms via programming to generate solutions, and in case the coding agent does not work, the reasoning agent will directly compute the results without programming. Extensive experiments on six graph analysis benchmarks demonstrate that GraphTeam achieves state-of-the-art performance with an average 25.85% improvement over the best baseline in terms of accuracy.
1 Introduction
@@
Figure 1: The overall pipeline of our multi-agent system GraphTeam (left), and the comparison between GraphTeam and the best baseline on six benchmarks (right).
Social networks Garton et al. (1997), urban computing Zheng et al. (2014), and various other fields make extensive use of graph data and related analytical methods. Given the promising problem-solving abilities demonstrated by large language models (LLMs) Brown et al. (2020); Anil et al. (2024); Anthropic (2024) across different domains, researchers are now exploring the feasibility of using LLMs to reason over graphs Tang et al. (2024); Sanford et al. (2024); Ouyang et al. (2024); Zhang et al. (2024a).
Existing LLM-based graph reasoning methods mainly fall into two categories. The first category Chai et al. (2023); Liu et al. (2023); Qin et al. (2023); Liu et al. (2024) usually combines LLMs with graph neural networks (GNNs) for modeling. A typical paradigm is to first encode the input graph with a GNN, and then feed the encodings to an LLM as language token embeddings. However, this line of work is specialized for graph machine learning tasks, especially node classification, and thus has limited transferability. The second category Fatemi et al. (2024); Chen et al. (2024c) usually flattens graphs into textual descriptions, and purely relies on LLMs for analyzing Liu and Wu (2025); Huang et al. (2023); Zhao et al. (2024); Shi et al. (2023); Chen et al. (2024c). Strategies like in-context learning Brown et al. (2020) and chain-of-thought reasoning Wei et al. (2022) are widely used in these methods to improve performance. However, these methods analyze graphs by step-by-step reasoning on raw inputs without utilizing any external knowledge or tools, severely limiting their performance.
Fortunately, LLMs have been shown to be capable of using external knowledge or tools for solving problems, known as a key feature of LLM-based agents Xi et al. (2023); Wang et al. (2024b). Moreover, inspired by the division of labor and cooperation mechanisms in human society, researchers propose LLM-based multi-agent systems Qian et al. (2024); Hong et al. (2024); Wu et al. (2024); Chen et al. (2024b), where agents with different specialities can collaborate with each other to address complex problems. But previous multi-agent frameworks are designed for general purposes, and have few knowledge relevant to graph reasoning.
@@
Inspired by human problem-solving processes Robson and Pólya (1945); Eisenberg and Berkowitz (1990); Bybee et al. (2006); Ramnarine-Rieks (2011), we present GraphTeam, a multi-agent system with three core modules specialized for graph reasoning: (1) Input-output normalization module. In human problem solving, clearly understanding and formalizing the problem is the first step Robson and Pólya (1945); Bybee et al. (2006), and answering the question in required formats is also important. Therefore, we introduce two different LLM-based agents: the question agent extracts and refines key arguments from the original question to help the system better understand the problem; the answer agent reformats the results to meet the specified output requirement. (2) External knowledge retrieval module. In tackling complex problems, humans often seek help from external sources Eisenberg and Berkowitz (1990) or draw analogies to previously solved ones Bybee et al. (2006). Therefore, we first build a knowledge base that includes relevant documentation and experience information. Then we introduce a search agent to retrieve the most relevant entries from this knowledge base to assist the downstream solving process. (3) Problem-solving module. Given the retrieved information, we introduce a coding agent that uses established algorithms via programming to generate solutions. If the coding agent fails, a reasoning agent directly infers the solution without relying on programming. The overall pipeline of GraphTeam is shown in Figure 1. We evaluate GraphTeam through extensive experiments across six graph reasoning benchmarks. These benchmarks cover a variety of graph tasks, ranging from basic graph theory to the deployment of GNNs. The experimental results and ablation studies highlight the effectiveness of GraphTeam as well as the necessity of each module. In summary, our contributions are three-fold:
@@
$\bullet$ Different from previous paradigms that utilize LLMs for graph reasoning, we propose to facilitate the solving process via the collaboration of multiple LLM-based agents.
@@
$\bullet$ Inspired by human problem-solving processes, we propose GraphTeam, a carefully designed framework consisting of five agents across three functional groups: input-output normalization, external knowledge retrieval, and problem-solving modules. The agents collaborate with a well-defined division of labor for graph reasoning.
@@
$\bullet$ Extensive experiments show that GraphTeam achieves state-of-the-art (SOTA) performance across all six graph reasoning benchmarks, with an average accuracy improvement of 15.3% over the best baseline.
2 Related Work
2.1 LLM-based Graph Reasoning
@@
To evaluate the ability of LLMs in solving graph reasoning problems, researchers have developed various benchmarks Guo et al. (2023); Fatemi et al. (2024); Das et al. (2024); Chen et al. (2024a); Li et al. (2024b). Among them, NLGraph Wang et al. (2024a) is one of the first benchmark for assessing LLMs’ graph analysis capabilities, focusing on basic graph theory reasoning and simple GNN computations. Graphwiz Chen et al. (2024a) evaluates LLMs on graph problems of varying complexity, including NP-complete ones. GraphInstruct Luo et al. (2024) includes 21 graph analysis tasks covering node-level, node-pair-level, and graph-level problems. Talk like a Graph Fatemi et al. (2024) measures basic graph understanding abilities, while LLM4DyG Zhang et al. (2024b) focuses on dynamic graph tasks related to temporal and spatial information. In our experiments, we employ the above five benchmarks, and also construct an additional one to evaluate the ability of deploying GNNs with AutoGL Guan et al. (2021).
@@
There is a rising trend of research to utilize LLMs for various graph reasoning tasks Li et al. (2024c); Jin et al. (2024). Here we focus on the most relevant methods that employ LLMs as predictors Chai et al. (2023); Liu et al. (2023); Qin et al. (2023); Liu et al. (2024). Related work can be roughly divided into two categories. The first category typically integrates LLMs with graph neural network (GNNs), and is specialized for graph machine learning tasks such as node classification Tang et al. (2024); Zhang et al. (2024a). For example, GraphGPT Tang et al. (2024) employs instruction tuning to fine-tune a GNN-enhanced LLM for node classification. GraphTranslator Zhang et al. (2024a) encodes nodes in a graph by a GNN, and then project node embeddings into the embedding space of LLM tokens for zero-shot classification. However, these methods are designed and optimized for a specific task, and cannot generalize well to the diverse reasoning tasks in the above benchmarks. The second category usually flatten graphs into textual descriptions of adjacency lists, and then rely on LLM itself to analyze graphs with strategies like in-context learning Brown et al. (2020), chain-of-thought reasoning Wei et al. (2022) for inference Liu and Wu (2025); Huang et al. (2023); Zhao et al. (2024); Shi et al. (2023); Chen et al. (2024c). For example, GUNDAM Ouyang et al. (2024) collects natural language-based question-answer pairs of graph analysis problems to fine-tune LLMs, and performs step-by-step reasoning for inference. However, recent studies have shown that the reasoning depths of current LLMs are still shallow Parmar et al. (2024); Li et al. (2024a), and will probably fail for problems that require complex reasoning. In contrast, we encourage LLMs to leverage external knowledge and tools, instead of performing step-by-step reasoning on raw inputs.
2.2 Multi-Agent Collaboration of LLMs
@@
Autonomous agents aim to solve problems with self-directed planning and actions. Recently, researchers introduce the concept of LLM-based agent that is capable of understanding natural language instruction, interacting with humans, perceiving external environments, and performing various actions Bang et al. (2023); Li et al. (2023); Xi et al. (2023); Wang et al. (2024b). Moreover, inspired by the collaboration of human in solving problems, researchers introduce LLM-based multi-agent systems, where agents collaborate with a well-defined division of labor. For example, ChatDev Qian et al. (2024) is a virtual software company where LLM-based agents are responsible for different software development tasks (such as user interface design or software testing) for collaboration. MetaGPT Hong et al. (2024) is a meta-programming framework that simulates human workflows. With proper developing pipeline, MetaGPT can generate effective programming solutions. AgentVerse Chen et al. (2024b) is a multi-agent framework inspired by human collaboration dynamics. It involves four iterative stages of expert recruitment, decision discussion, action and evaluation. AutoGen Wu et al. (2024) is another multi-agent framework for general purposes. Through tool invocation and code generation, AutoGen can accomplish a wide variety of tasks. MACRec Wang et al. (2024c) is a novel framework designed to enhance recommendation systems through multi-agent collaboration. However, these multi-agent frameworks are not specialized for graph reasoning, e.g., they have few knowledge of relevant Python libraries and no experience base of similar problems.
3 Methodology
@@
We start by presenting the overall framework of GraphTeam, and then present the details of each agent in the system.
3.1 Framework Overview
@@
As shown in Figure 2, the entire framework of GraphTeam consists of five agents across three functional groups:
Figure 2: The overall framework of GraphTeam, which includes five agents across three functional groups. Firstly, the question agent refines the original question by extracting key arguments. Then the search agent queries knowledge bases with the reformatted question to retrieve relevant experience and documentation. Subsequently, the coding agent attempts to solve the problem via programming. If the coding agent fails after several retries, the reasoning agent will take over and directly solve the question without programming. Finally, the solutions from either the coding or reasoning agents are then passed to the answer agent for format standardization.
@@
External Knowledge Retrieval: This functional group only includes the search agent to extract relevant information from external knowledge base. Specifically, we first build a knowledge base for graph reasoning based on documentation of popular Python libraries and experiences of previous problem-solving processes. With problems refined by the upstream question agent as queries, the search agent then retrieves information beneficial to downstream problem solving from the knowledge base.
Problem Solving: This functional group calculates the results based on the refined question of question agent and the extracted knowledge of search agent. Specifically, the coding agent attempts to solve the problem via programming. If the coding agent fails after several retries, the reasoning agent will solve the problem directly without programming.
3.2 Question Agent
Abstraction is a preliminary step for human beings to solve problems Robson and Pólya (1945). Inspired by this, we introduce a question agent to extract and refine four key arguments from the original problem descriptions. Formally, given a graph reasoning problem with its original description denoted as $q$, we employ an LLM denoted as $\operatorname{Question}(q)$ to generate refined question $q_{r}$, graph type $q_{t}$, input graph $q_{g}$ and output format $q_{f}$:
$(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q),$ (1)
where $q_{r}$ is a condensed version of original instruction, $q_{t}$ indicates the graph type (e.g., directed, undirected, weighted, etc.), $q_{g}$ contains the graph to be processed, and $q_{f}$ specifies how the answer should be presented (e.g., numerical value, list, etc.). By explicitly identifying these key information, downstream agents can better understand a task and its requirements, thereby enhancing the overall reasoning capability. An example of question agent is shown below: